From pmw@roe.ac.uk Mon Jun 9 14:49:48 2008 Date: Mon, 9 Jun 2008 13:54:16 +0100 From: Peredur Williams To: Peredur Williams , Nigel Hambly , Eckhard Sutorius , Mike Read , Mark Holliman , Nicholas Cross , Rob Blake , Ross Collins Cc: Andy Lawrence , Andy Adamson , Brian Walshe , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 9/06/08 Minutes of WFAU VDFS Science Archive meeting: 9th June 2008 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: RSC, NJC, MAR, LGR, MSH, RPB, PMW Apologies: NCH, ETWS, JPE, JMS, AL, BCW, RGM DoNM: 10am, Friday 13th June 2008 in the Vista Hut (yup, this week) Actions discharged this week: ----------------------------- Actions partly discharged but continuing: ----------------------------------------- ACTION: NJC to test CU13/14 from newly merged SVN trunk in advance of running by ops for DXS in DR4 Python code works. Asked CASU for for bugfix or use old version of CFITSIO. ACTION: ETWS to run CU4 for individual UKIDSS surveys after successful backup of the ingest DB Underway. ACTION: MAR (with advise from NCH) to progress as much as possible of QC2 Underway. ACTION: ETWS to continue with routine CUs 1-3 to ensure timely ingest & release of 08A processed flat files for Mar/Apr etc. Underway. ACTION: RPB to formulate backup plan arising from meeting above Continues. Project management: NJC attended the VIDEO meeting in Hertforshire on June 4th and has placed a report on the Twiki at: http://apache.roe.ac.uk/twiki/bin/view/WFAU/VIDEOSurveyMeeting WFCAM & VISTA updates: Nothing for a while. Comments and issues arising from CASU fortnightly minutes: No new minutes this week. Networking: WSA Operations: ETWS reported (by e-mail) that the first run of the CU4 recalibration had finished, out of 810 days there were problems on 120 days where the data hadn't been fully ingested. This was mainly due either to a missing extension and a problem in writing default entries for this extension or to observations outside of the illumination correction table positions which weren't caught correctly in the code. He had fixed both bugs in the code and re-processed this data which is now as well ingested. There are a few different other problems left due to singular files, but in general he hoped the recalibration would be completed latest on Wednesday. Hardware MSH reported revised plan for Thoth: he was taking it over; the motherboard would be replaced tomorrow (Tuesday) as planned, but the upgrade postponed. Installation of 64-bit Debian 4 and software on Horus causing problems, glibc 2.5 not compiling, may have to use 32-bit Starlink software. Software: RSC reported that he had tested out modifications to CU19 for a split two-phase data release that would exclude the GPS in the first part, as the software was a bit messy at doing this for the last release we expect to be doing the same for this release. In lieu of bugs to fix, he'd started investigating the rewrite of CU7 to make use of the MergeSource table to minimise the impact of recalibrations, which would take some time. ETWS reported that he had also worked on the installation of the required third-party software on khafre, nearly everything compiled fine with the existing gcc-4.1.2, so khafre might be back in service at the end of the week. Survey Data Release: Because of hiccups with reingestion of recalibrated data it was decided to revisit the schedule at the next meeting, this Friday, when the position should be clearer Non-survey Data Release: Astrogrid deployment & Data Analysis services Miscellaneous: Nothing else this week.